OpenStudy (kj4uts):

Suppose an isosceles triangle ABC has A=pi/4 and b=c=4. What is the length of a^2? Please explain. Thank you!

OpenStudy (trojanpoem):



OpenStudy (trojanpoem):

a^2 = (4)^2+(4)^2 - (2 * 4 * 4 * cos(45))

OpenStudy (trojanpoem):

this is an iso so <ACB = ABC = 180 - 45/ 2 = 135 /2 = 67.5 so 4/sin(67.5) = a / sin45 a = 4*sin45/sin(67.5)

OpenStudy (trojanpoem):

If you want more solutions contact me.